Mir Status Report 5 October 1997

NASA MIR Program
Mir 24 / NASA 6 Status

Date: Sunday, October 5, 1997
Mission Day: Mir 24 / NASA 6 MD 61/10

Mir 24 CDR Anatoly Solovyov
Mir 24 FE Pavel Vinogradov
NASA 6 FE Dave Wolf

Mir Station System/Subsystem Activities
In preparation for tomorrow's undocking, the crew closed the hatch to
Progress and performed leak checks today and also performed a TORU checkout.
Progress M-36 launched successfully early this evening.

Progress Information
Progress M-36 launch: Sunday 18:08 DMT
Progress M-35 undock: Monday 14:24 DMT
Progress M-37 dock: Tuesday 19:41 DMT

Mir Payload Activities
HLS:MS:I:8 Back Pain Questionnaire
MG:FAC:I:39 BEM/BSTC-M & BTR Facility Maintenance
SMP:CHS:I:2 Heart Watch Monitoring of Exercise

Nothing to report.

Data Management
Nothing to report.

Network Status
Nothing to report.

Ground Facilities Status
Nothing to report.

Nothing to report.

Experiment Operations
Dave reported that the MIM flotor is doing great.

Experiment Hardware
Dave requested instructions on how to install Russian fonts on COSS so that he can read Russian radiograms on that machine.

Dave reported that the right-hand video unit on CAPE did not appear to be taking data. Dave changed the interval time from one hour to 30 minutes and the video unit took a picture immediately. However, when he changed the setting back to one hour, no action resulted. POSA is working with the PIs to draft instructions on this.

Nothing to report.

Nothing to report.

Dave's first Mir PFC was completed today. Backpain questionnaire was
scheduled for today.

Environmental Parameters (previous 24 hour period):
Parameter Min. Max. Nominal Range
Pressure (mmHg) Base Block 747 753 660-850
ppO2 (mmHg) 175 180 140-200
ppCO2 (mmHg) 3.5 4.2 up to 8.0
ppH20 (mmHg) 12.4 13.2 8-14
Humidity (percentage) 50 53 30-75
Temperature (xC) base block (BB) 24.8 26.0 18-28

Communications Summary
A/G with Consultant Group
10:42-10:47 COSS Russian fonts
17:10-17:18 CAPE, COSS
19:55-20:05 PFC

Packet --
Three work files (form 24, earth obs and COSS instructions) and five personal files were submitted for uplink.

BIO-3D operations are still in question.


MIR 24/NASA 6 MD 62/11
HLS:MS:I:8 Back Pain Questionnaire
MG:ENV:I:7 SAMS undock
ISS:DYN:I:80 MiSDE undock
CGBA I:17 Status check
BIO3D I:39 Status check
BIO3D Hardware setup
BIO3D I:60 Cult/Photo
SMP:MO:I:44 TEPC Calldown

Other Activities
Progress undocking

  • submit to reddit


Support SpaceRef, NASA Watch and the Astrobiology Web on Patreon.

Join our mailing list

Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on October 5, 1997 8:36 PM.

Mir Status Report 24 September 1997 was the previous entry in this blog.

Mir Status 19 October 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.